# Hadronic molecules with four charm or beauty quarks

### v1

@@ Abstract
We apply the extended local hidden gauge formalism to study the meson-meson interactions with the quark constituents $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$, where the exchanged mesons can only be the fully-heavy vector mesons $J/\psi$, $B_{c}^{*}$, and $\Upsilon$. We solve the coupled-channel Bethe-Salpeter equation to derive two poles in the $bb\bar{c}\bar{b}$ system and two poles in the $cc\bar{c}\bar{b}$ system (with four charge-conjugated poles in the $\bar{b}\bar{b}cb$ and $\bar{c}\bar{c}cb$ systems): one pole existing in the $bb\bar{c}\bar{b}$ system corresponds to the sub-threshold bound state when setting the cutoff momentum Λ>850 $\Lambda>850\textrm{ MeV}$, while it causes the threshold effect with Λ<850 $\Lambda<850\textrm{ MeV}$; the other pole existing in this system corresponds to the sub-threshold bound state with Λ>1100 $\Lambda>1100\textrm{ MeV}$; the two poles existing in the $cc\bar{c}\bar{b}$ system correspond to the sub-threshold bound states only when setting Λ>1550 $\Lambda>1550\textrm{ MeV}$ and Λ>2650 $\Lambda>2650\textrm{ MeV}$, making them difficult to be identified as deeply-bound hadronic molecules. We propose to investigate the two poles of the $bb\bar{c}\bar{b}$ system in the $\mu^{+}\mu^{-}B_{c}^{-}$ channel at LHC.
@@ I Introduction
In recent decades, the study of exotic hadrons has gradually become a focal point in hadron physics. Some exotic hadrons exhibit multiquark compositions, such as the compact tetraquark states and the meson-meson molecular states. The picture of hadronic molecules has achieved significant success in the light-quark sector Oller:1997ti ; Oller:1998hw ; Oset:1997it ; Jido:2003cb ; Bruns:2010sv ; Garcia-Recio:2003ejq ; Hyodo:2002pk , which can be used to explain many resonances, such as the $f_{0}(980)$ and $a_{0}(980)$, etc. Besides, many hidden-charm pentaquark states observed in the past decade can be interpreted as the hadronic molecules that are dynamically generated through the meson-baryon interactions within the local hidden gauge framework Wu:2010jy ; Wu:2010vk ; Chen:2015sxa ; He:2015cea ; Xiao:2013yca ; Roca:2015dva ; Liu:2015fea ; Uchino:2015uha . In recent years, several exotic structures in the di-$J/\psi$ invariant mass spectrum have been reported by the LHCb, CMS, and ATLAS collaborations LHCb:2020bwg ; ATLAS:2023bft ; CMS:2023owd , including the $X(6200)$, $X(6600)$, $X(6900)$, and $X(7200)$. These structures are good candidates for the fully-charmed tetraquark states. Extensive theoretical investigations have been performed to elucidate their nature liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Bedolla:2019zwg ; Weng:2020jao ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Zhao:2020zjh ; Mutuk:2021hmi ; Wang:2021kfv ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Feng:2020riv ; Ma:2020kwb ; Maciula:2020wri ; Goncalves:2021ytq ; Wang:2020gmd ; Esposito:2021ptx ; Zhuang:2021pci ; Zhao:2020nwy ; Becchi:2020uvq ; Sonnenschein:2020nwn ; Zhu:2020snb ; Wan:2020fsk ; Gordillo:2020sgc ; Liu:2020tqy ; Majarshin:2021hex ; Kuang:2022vdy ; Wang:2021mma ; Chen:2016jxd ; Czarnecki:2017vco , some of which attempted to explore their nature as molecular states Guo:2020pvt ; Cao:2020gul ; Gong:2020bmg ; Dong:2021lkh ; Ortega:2023pmr ; Wang:2022yes ; Zhou:2022xpd , but a definitive and conclusive resolution remains elusive.
@@
Previous theoretical studies on the fully-heavy tetraquark states mainly focus on the interpretation of compact tetraquark states, while there are not so many studies based on the interpretation of hadronic molecular states. This is because the exchanged hadrons of these systems all have quite large masses, such as the fully-heavy vector mesons $J/\psi$, $B_{c}^{*}$, and $\Upsilon$ exchanged in the $cb\bar{c}\bar{b}$ system within the extended local hidden gauge formalism, so their induced interactions are significantly suppressed. In Ref. Liu:2023gla we studied the $cb\bar{c}\bar{b}$ system to explore the existence of the fully-heavy hadronic molecules $B_{c}^{(*)}\bar{B}_{c}^{(*)}$. Within the extended local hidden gauge framework, we found that the two fully-heavy mesons $B_{c}^{(*)}$ and $\bar{B}_{c}^{(*)}$ are able to form a bound state by exchanging the relatively lighter meson $J/\psi$.
@@
In this paper we apply the extended local hidden gauge formalism to further investigate the fully-heavy hadronic molecules existing in the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. By solving the coupled-channel Bethe-Salpeter equation, we evaluate the hadronic molecules generated from the meson-meson interactions in these systems. Our results suggest the existence of two potential bound states in the $bb\bar{c}\bar{b}$ system (with two charge-conjugated states in the $\bar{b}\bar{b}cb$ system), while they manifest as the threshold effects as the cutoff momentum varies. Both of these two structures share the same spin-parity quantum number $J^{P}=1^{+}$, and their possible observing channel is $\mu^{+}\mu^{-}B_{c}^{-}$ at LHC. Besides, we find two poles in the $cc\bar{c}\bar{b}$ system (with two charge-conjugated poles in the $\bar{c}\bar{c}cb$ system), but it is difficult to identify them as deeply-bound hadronic molecules.
@@
In Ref. Liu:2023gla we have applied the extended local hidden gauge formalism to study the interactions of the $cb\bar{c}\bar{b}$ system. In this section we follow the same approach to study the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. Within this framework, the interactions are primarily contributed by the vector meson exchange, as depicted in Fig. 1. The corresponding Lagrangians are written as:
@@
The coupling constant $g$ is generally defined as $g={M_{V}}/({2f_{P}})$, with $M_{V}$ the mass of the exchanged vector meson and $f_{P}$ the decay constant of its corresponding pseudoscalar meson. Since the charm and bottom quarks do not form a flavor $SU(2)$ symmetry, we can not use an overall parameter any more. For the exchange of the $J/\psi$, $B_{c}^{*}$, and $\Upsilon$ mesons, we respectively use
@@
where $p_{1}$ and $p_{2}$ are four-momenta of the incoming mesons; $p_{3}$ and $p_{4}$ are four-momenta of the outgoing mesons; $\epsilon_{1}$ and $\epsilon_{2}$ are polarization vectors of the incoming mesons; $\epsilon_{3}$ and $\epsilon_{4}$ are polarization vectors of the outgoing mesons. We use the subscripts $PP$, $VP$, and $VV$ to denote the pseudoscalar-pseudoscalar, vector-pseudoscalar, and vector-vector sectors, respectively. We use the superscripts $t$, $u$, and $co$ to denote the vector meson exchange in the $t$ and $u$ channels as well as the contact term, respectively.
@@ II.1 The $cc\bar{c}\bar{c}$ system
In the $cc\bar{c}\bar{c}$ system, the $PP$ interaction involves only one channel $\eta_{c}\eta_{c}$, the $VP$ interaction involves only one channel $J/\psi\eta_{c}$, and the $VV$ interaction also involves only one channel $J/\psi J/\psi$.
@@
$C_{PP}^{t}$ $=$ $0\,,$
$C_{PP}^{u}$ $=$ $0\,,$
$C_{VP}^{t}$ $=$ $0\,,$ (11)
$C_{VV}^{t}$ $=$ $0\,,$
$C_{VV}^{u}$ $=$ $0\,,$
$C_{VV}^{co}$ $=$ $0\,.$
Therefore, our results do not support the existence of hadronic molecules in the $cc\bar{c}\bar{c}$ system.
@@ II.2 The $cc\bar{c}\bar{b}$ and $\bar{c}\bar{c}cb$ systems
The results for the $cc\bar{c}\bar{b}$ and $\bar{c}\bar{c}cb$ systems are the same, so we only need to study the $cc\bar{c}\bar{b}$ system. In this system, the $PP$ interaction involves only one channel $\eta_{c}B_{c}^{+}$, whose coefficients are
@@
The reduction factor $\lambda$ is introduced to account for the large mass difference between the initial and final mesons. Following Ref. Yu:2018yxl , numerically we use
@@
The coefficient $C_{PP}^{u}$ is positive, indicating the interaction due to the exchange of the $B_{c}^{*}$ meson to be repulsive, so there do not exist hadronic molecules in the $PP$ sector.
The $VP$ interaction involves two coupled channels $J/\psi B_{c}^{+}$ and $B_{c}^{*+}\eta_{c}$, whose coefficient is
$C_{VP}^{t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&J/\psi B_{c}^{+}&B_{c}^{*+}\eta_{c}\\ \hline\cr J/\psi B_{c}^{+}&0&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\\ B_{c}^{*+}\eta_{c}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\end{array}\right)\,.$ (17)
@@
$C_{VV}^{t}$ $=$ $0\,,$ (24)
$C_{VV}^{u}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,,$
and the relevant contact term is

### v2

@@ Abstract
We apply the extended local hidden gauge formalism to study the meson-meson interactions with the quark constituents $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$, where the exchanged mesons are the fully-heavy vector mesons $J/\psi$, $B_{c}^{*}$, and $\Upsilon$. We solve the coupled-channel Bethe-Salpeter equation to derive two poles in the $bb\bar{c}\bar{b}$ system and two poles in the $cc\bar{c}\bar{b}$ system. There are also four charge-conjugated poles in the $\bar{b}\bar{b}cb$ and $\bar{c}\bar{c}cb$ systems. In the $bb\bar{c}\bar{b}$ system, one pole corresponds to a sub-threshold bound state when the cutoff momentum is set to Λ>850 $\Lambda>850\textrm{ MeV}$. The other pole in this system corresponds to a sub-threshold bound state when Λ>1100 $\Lambda>1100\textrm{ MeV}$. In the $cc\bar{c}\bar{b}$ system, the two poles correspond to sub-threshold bound states only when Λ>1550 $\Lambda>1550\textrm{ MeV}$ and Λ>2650 $\Lambda>2650\textrm{ MeV}$. This makes them difficult to identify as deeply-bound hadronic molecules. We propose investigating the two poles of the $bb\bar{c}\bar{b}$ system in the $\mu^{+}\mu^{-}B_{c}^{-}$channel at the LHC.
@@ I Introduction
In recent decades, the study of exotic hadrons has gradually become a focal point in hadron physics. Some exotic hadrons exhibit multiquark compositions, such as the compact tetraquark states and the meson-meson molecular states Liu:2019zoy ; Chen:2022asf ; Guo:2017jvc ; Brambilla:2019esw ; Esposito:2016noz ; Lebed:2016hpi ; Ali:2017jda . The picture of hadronic molecules has achieved significant success in the light-quark sector Oller:1997ti ; Oller:1998hw ; Oset:1997it ; Jido:2003cb ; Bruns:2010sv ; Garcia-Recio:2003ejq ; Hyodo:2002pk , which can be used to explain many resonances, such as the $f_{0}(980)$ and $a_{0}(980)$, etc. Besides, many hidden-charm pentaquark states observed in the past decade can be interpreted as the hadronic molecules that are dynamically generated through the meson-baryon interactions within the local hidden gauge framework Wu:2010jy ; Wu:2010vk ; Chen:2015sxa ; He:2015cea ; Xiao:2013yca ; Roca:2015dva ; Liu:2015fea ; Uchino:2015uha . In recent years, several exotic structures in the di-$J/\psi$ invariant mass spectrum have been reported by the LHCb, CMS, and ATLAS collaborations LHCb:2020bwg ; ATLAS:2023bft ; CMS:2023owd , including the $X(6200)$, $X(6600)$, $X(6900)$, and $X(7200)$. These structures are good candidates for the fully-charmed tetraquark states. Extensive theoretical investigations have been performed to elucidate their nature liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Bedolla:2019zwg ; Weng:2020jao ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Zhao:2020zjh ; Mutuk:2021hmi ; Wang:2021kfv ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Feng:2020riv ; Ma:2020kwb ; Maciula:2020wri ; Goncalves:2021ytq ; Wang:2020gmd ; Esposito:2021ptx ; Zhuang:2021pci ; Zhao:2020nwy ; Becchi:2020uvq ; Sonnenschein:2020nwn ; Zhu:2020snb ; Wan:2020fsk ; Gordillo:2020sgc ; Liu:2020tqy ; Majarshin:2021hex ; Kuang:2022vdy ; Wang:2021mma ; Chen:2016jxd ; Czarnecki:2017vco , some of which attempted to explore their nature as molecular states Guo:2020pvt ; Cao:2020gul ; Gong:2020bmg ; Dong:2021lkh ; Ortega:2023pmr ; Wang:2022yes ; Zhou:2022xpd , but a definitive and conclusive understanding of their nature remains elusive.
@@
Previous theoretical studies on the fully-heavy tetraquark states mainly focus on the interpretation of compact tetraquark states, while there are not so many studies based on the interpretation of hadronic molecular states. This is because the exchanged hadrons of these systems have quite large masses, such as the fully-heavy vector mesons $J/\psi$, $B_{c}^{*}$, and $\Upsilon$ exchanged in the $cb\bar{c}\bar{b}$ system within the extended local hidden gauge formalism, so their induced interactions are significantly suppressed. In Ref. Liu:2023gla we studied the $cb\bar{c}\bar{b}$ system to explore the existence of the fully-heavy hadronic molecules $B_{c}^{(*)}\bar{B}_{c}^{(*)}$. Within the extended local hidden gauge framework, we found that the two fully-heavy mesons $B_{c}^{(*)}$ and $\bar{B}_{c}^{(*)}$ are able to form a bound state by exchanging the relatively lighter meson $J/\psi$.
@@
In this paper we apply the extended local hidden gauge formalism to further investigate the fully-heavy hadronic molecules existing in the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. By solving the coupled-channel Bethe-Salpeter equation, we evaluate the hadronic molecules generated from the meson-meson interactions in these systems. Our results indicate the possible existence of two bound states in the $bb\bar{c}\bar{b}$ system, along with two charge-conjugated states in the $\bar{b}\bar{b}cb$ system. However, their manifestation depends on the cutoff momentum, where they may appear as threshold effects. Both structures share the same spin-parity quantum number $J^{P}=1^{+}$, and can potentially be observed in the $\mu^{+}\mu^{-}B_{c}^{-}$ channel at LHC. Besides, we find two poles in the $cc\bar{c}\bar{b}$ system (with two charge-conjugated poles in the $\bar{c}\bar{c}cb$ system), but it is difficult to identify them as deeply-bound hadronic molecules.
@@
In Ref. Liu:2023gla we have applied the extended local hidden gauge formalism to study the interactions of the $cb\bar{c}\bar{b}$ system. In this section we follow the same approach to study the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. Note that the validity of the extended local hidden gauge formalism in these systems is still questionable. Within the framework of local hidden gauge  Bando:1987br ; Meissner:1987ge , the vector mesons are considered to act as gauge bosons, transmitting interactions. This mechanism successfully describes many low-energy interactions Wu:2010jy ; Oset:2010tof ; Aceti:2014uea ; Geng:2008gx ; Nagahiro:2008cv , especially in processes dominated by the exchange of light vector mesons. Unfortunately, it is not clear whether this mechanism remains effective in the heavy flavor region. An extension of this method has also been developed to study molecular state candidates containing heavy quarks Wu:2010jy ; Roca:2015dva ; Molina:2009ct . In these cases, generally, the exchange of light vector mesons and heavy vector mesons is allowed simultaneously. Taking the vector meson-vector meson interaction in the $c\bar{c}s\bar{s}$ system as an example, this system involves two coupling channels: $D_{s}^{*}\bar{D}_{s}^{*}$ and $J/\psi\phi$. In the $D_{s}^{*}\bar{D}_{s}^{*}$ channel, the exchange of $\phi$ and $J/\psi$ occurs simultaneously. Compared to the $\phi$ exchange, the $J/\psi$ exchange is considered to be secondary due to its larger mass. However, things are somewhat different in the $J/\psi\phi$ channel. The minimal constituents for the exchanged meson should be $c\bar{c}s\bar{s}$, which we do not consider in the local hidden gauge formalism. Such type of exchange mechanism should also be suppressed due to their involvement in the exchange of four quarks. On the other hand, $J/\psi\phi$ channel can couple to $D_{s}^{*}\bar{D}_{s}^{*}$ channel, which allows for the exchange of the heavy vector meson $D_{s}^{*}$ between these two channels. Although the exchange of heavy vector mesons should be relatively small, it plays a role in this scenario. Under this assumption, results consistent with experiments can be obtained Molina:2009ct . So the present study as well as Ref. Liu:2023gla serve as pioneer researches to investigate the fully-heavy hadronic molecules. Within this framework, the interactions are primarily contributed by the vector meson exchange, as depicted in Fig. 1. The corresponding Lagrangians are written as:
@@
The coupling constant $g$ is generally defined as $g={M_{V}}/({2f_{P}})$, with $M_{V}$ the mass of the exchanged vector meson and $f_{P}$ the decay constant of its corresponding pseudoscalar meson. Since the charm and bottom quarks do not form a flavor $SU(2)$ symmetry, we can not use an overall parameter. For the exchange of the $J/\psi$, $B_{c}^{*}$, and $\Upsilon$ mesons, we respectively use
@@
where $p_{1}$ and $p_{2}$ are four-momenta of the incoming mesons; $p_{3}$ and $p_{4}$ are four-momenta of the outgoing mesons; $\epsilon_{1}$ and $\epsilon_{2}$ are polarization vectors of the incoming mesons; $\epsilon_{3}$ and $\epsilon_{4}$ are polarization vectors of the outgoing mesons. We use the subscripts $PP$, $VP$, and $VV$ to denote the pseudoscalar-pseudoscalar, vector-pseudoscalar, and vector-vector sectors, respectively. We use the superscripts $t$, $u$, and $co$ to denote the vector meson exchange in the $t$ and $u$ channels as well as the contact term, respectively. Note that the coefficient $C_{VP}^{u}$ is zero because we only consider the vector meson exchange in the present study. Actually, there also exists the pseudoscalar meson exchange. As detailedly evaluated in Ref. Aceti:2014kja ; Oset:2022xji , the contribution of the pseudoscalar meson exchange is negligible near the threshold, compared to the vector meson exchange. This is because the amplitude of the vector meson exchange is proportional to the energy of the external meson, while the amplitude of the pseudoscalar meson exchange is proportional to the three-momentum of the external meson that is always negligible near the threshold.
@@ II.1 The $cc\bar{c}\bar{c}$ system
In the $cc\bar{c}\bar{c}$ system, the $PP$ interaction involves only one channel $\eta_{c}\eta_{c}$, the $VP$ interaction involves only one channel $J/\psi\eta_{c}$, and the $VV$ interaction also involves only one channel $J/\psi J/\psi$. Besides, the $bb\bar{b}\bar{b}$ system can be similarly investigated.
@@
$C_{PP}^{t}$ $=C_{PP}^{u}=0\,,$
$C_{VP}^{t}$ $=C_{VP}^{u}=0\,,$
$C_{VV}^{t}$ $=C_{VV}^{u}=0\,,$
$C_{VV}^{co}$ $=0\,.$ (11)
The $bb\bar{b}\bar{b}$ system and the $cc\bar{c}\bar{c}$ system exhibit very similar dynamic properties, with the only distinction being their mass differences. Consequently, they share the same coefficients listed in Eq. II.1, implying that the interactions within the $bb\bar{b}\bar{b}$ system also vanish.
Therefore, our results do not support the existence of hadronic molecules in the $cc\bar{c}\bar{c}$ and $bb\bar{b}\bar{b}$ systems. Moreover, our conclusions actually do not depend on the value of the coupling constant $g$. As long as the near-threshold interactions in these systems are dominated by vector meson exchange and can be described by the Lagrangian in Eq. 1, the validity of the conclusions can be maintained. Let’s examine the specific form of the $t$-channel exchange potential in Eq. II,
$-iV^{t}=ig_{1}(p_{1}+p_{3})^{\mu}ig_{2}(p_{2}+p_{4})^{\nu}\frac{i(-g_{\mu\nu}+\frac{q_{\mu}q_{\nu}}{m_{V}^{2}})}{t-m_{V}^{2}}\,,$ (12)
where $p_{1}$ and $p_{2}$ are four-momenta of the incoming mesons; $p_{3}$ and $p_{4}$ are four-momenta of the outgoing mesons; $q$ is four-momenta of the exchanged meson and $g_{1}$ and $g_{2}$ are two coupling constants corresponding to top and bottom vertex in the fig. 1 (a). Ignoring the three-momenta of the particles and approximating $t$ as 0, the following simplification can be obtained,
$V^{t}\simeq{{g_{1}g_{2}}\over{m_{V}^{2}}}(m_{1}+m_{3})\cdot(m_{2}+m_{4}).$ (13)
It can be found that if the particles in the channels are identical, i.e., $g_{1}=g_{2}$, the interaction will be repulsive. Similar derivations also apply to the exchange potential $V^{u}$ in the $u$-channel. This indicates that the existence of near-threshold molecular states in channels similar to $\eta_{c}\eta_{c}$ is difficult to support. The experimentally observed states in these systems are therefore good candidates to be tetraquark states.
@@ II.2 The $cc\bar{c}\bar{b}$ and $\bar{c}\bar{c}cb$ systems
The results for the $cc\bar{c}\bar{b}$ and $\bar{c}\bar{c}cb$ systems are the same, so we only need to study the $cc\bar{c}\bar{b}$ system. Besides, the $bb\bar{c}\bar{b}$ and $\bar{b}\bar{b}cb$ systems can be similarly investigated. In the $cc\bar{c}\bar{b}$ system, the $PP$ interaction involves only one channel $\eta_{c}B_{c}^{+}$, whose coefficients are
@@
The reduction factor $\lambda$ existing in the $u$ channel is introduced to account for the large mass difference between the initial meson $\eta_{c}$ and the final meson $B_{c}^{+}$ (or between the initial meson $B_{c}^{+}$ and the final meson $\eta_{c}$). Following Ref. Yu:2018yxl , numerically we use
@@
The coefficient $C_{PP}^{u}$ is positive, indicating the interaction due to the exchange of the $B_{c}^{*}$ meson is repulsive, hence hadronic molecules in the PP sector are not expected to exist.
The $VP$ interaction involves two coupled channels $J/\psi B_{c}^{+}$ and $B_{c}^{*+}\eta_{c}$, whose coefficient are
$C_{VP}^{t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&J/\psi B_{c}^{+}&B_{c}^{*+}\eta_{c}\\ \hline\cr J/\psi B_{c}^{+}&0&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\\ B_{c}^{*+}\eta_{c}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\end{array}\right)\,\quad and$ (19)
$C_{VP}^{u}$ $=$ ${\bf 0}_{2\times 2}\,.$
@@
$C_{VV}^{t}$ $=$ $0\,\quad and$ (26)
$C_{VV}^{u}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,.$
The relevant contact term is